OpenStudy (3psilon):

I could never get these types of questions, please help! From a jar containing 50 pieces of candy , of which 25 are red and 25 are green, Ari has taken 3 red and 4 green pieces . He takes an additional 13 pieces from the jar. What is the least number of these additional pieces that must be red in order for Ari to have more red candies than green candies among all the pieces he has taken

OpenStudy (compassionate):

r = 25 g = 25 - 3r & - 4g ---------- r = 22 I - 13 I g = 21 I

OpenStudy (3psilon):

@Compassionate I already got that far. Just confused with the rest